OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 2r2 + 3s3 - r2 + 4t2 - r2 if r = -2, s = -3, and t = 5. -19 -37 27 19

OpenStudy (anonymous):

@kewlgeek555 @helpmeplzzzzz!

OpenStudy (anonymous):

does 2r2 = \[2r^{2}\] ?

OpenStudy (anonymous):

Your question is even easier thank your username.. :P

OpenStudy (anonymous):

Yes, sorry it is Evaluate the following expression using the values given: Find 2r^2 + 3s^3 - r^2 + 4t^2 - r^2 if r = -2, s = -3, and t = 5. -19 -37 27 19

OpenStudy (anonymous):

2*4 + 3* (-27) + 4*25 -4 = 8-81+100 -4 = 23

OpenStudy (anonymous):

Thanks @Allofus

OpenStudy (anonymous):

But that isn't an option

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

I know the answer isn't 27 either

OpenStudy (anonymous):

HELP PEOPLES

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@waterineyes @helpmeplzzzzz! @Hotchellerae21

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Paying attention is the hard part. 2r^2 + 3s^3 - r^2 + 4t^2 - r^2 r = -2 2(-2)^2 + 3s^3 - (-2)^2 + 4t^2 - (-2)^2 2*4 + 3s^3 - 4 + 4t^2 - 4 8 + 3s^3 + 4t^2 - 8 3s^3 + 4t^2 s = -3 3(-3)^3 + 4t^2 3*(-27) + 4t^2 -81 + 4t^2 t = 5 -81 + 4(5)^2 -81 + 4*25 -81 + 100 19 Keep in mind that -2^2 and (-2)^2 are unlikely to be the same in written notation. They may be in programming languages.

OpenStudy (idealist10):

2(-2)^2+3(-3)^3-(-2)^2+4(5)^2-(-2)^2 =2(4)+3(-27)-(4)+4(25)-(4) =8-81-4+100-4 =-73-4+100-4 =-77+100-4 =23-4 =19 Answer: 19

OpenStudy (anonymous):

\[2 r^2-r^2-r^2+3 s^3+4 t^2=3 s^3+4 t^2 \]\[3 (-3)^3+4\ 5^2=19 \]